Compute the plot and probability distribution for a sample of 5 items taken at random from a production process known to produce 30% defective items.
Answer by Boreal(15235)

probability of 0 is 0.7^5=0.1681
for 1 it is 5*0.3*0.7^4=0.3602
for 2 it is 10*0.3^2*0.7^3=0.3087
for 3 it is 10*0.3^3*0.7^2=0.1323
for 4 it is 5*0.3^4*0.7=0.0284
for 5 it is 0.3^5=0.00243